# The strong Fe K line and spin of the black-hole X-ray binary MAXI J1631–479

### v1

@@ Abstract
We study the transient black hole binary MAXI J1631–479 in its soft spectral state observed simultaneously by the NICER and NuSTAR instruments. Its puzzling feature is the presence of a strong and broad Fe K line, while the continuum consists of a strong disk blackbody and a very weak power-law tail. The irradiation of the disk by a power-law spectrum fitting the tail is much too weak to account for the strong line. Two solutions were proposed in the past. One invoked an intrinsic Fe K disk emission, and the other invoked disk irradiation by the returning blackbody emission. We instead find that the strong line is naturally explained by the irradiation of the disk by the spectrum from Comptonization of the disk blackbody by coronal relativistic electrons. The shape of the irradiating spectrum at $\lesssim$10 keV reflects that of the disk blackbody; it is strongly curved and has a higher flux than that of a fit with a power-law irradiation. That flux accounts for the line. While this result is independent of the physical model used for the disk intrinsic emission, the value of the fitted spin strongly depends on it. When using a Kerr disk model for a thin disk with a color correction, the fitted spin corresponds to a retrograde disk, unlikely for a Roche-lobe overflow binary. Then, a model accounting for both the disk finite thickness and radiative transfer yields a spin of $a_{*}\approx 0.8$–0.9, which underlines the strong model-dependence of X-ray spin measurements.
@@ 1 Introduction
Its BH mass, $M$, and the distance, $D$, remain uncertain. Rout et al. (2023) claimed a very high BH mass, $26{M}_{\sun}$, with a preference for $\sim\!60{M}_{\sun}$. This followed from a disk blackbody model at a high inclination, $i\approx 71\pm 1\arcdeg$, found by them. On the other hand, low inclination values were found in the spectral studies by Xu et al. (2020), $i=29\pm 3\arcdeg$, and by Draghis et al. (2024), $i=22^{+10}_{-12}\arcdeg$. Its dimensionless spin parameter was estimated to be very high, $a_{*}>0.94$ by Xu et al. (2020), and $a_{*}>0.996$ by Rout et al. (2023). Both studies used the X-ray reflection method (Bambi et al., 2021) in the soft spectral states of MAXI J1631–479. However, the results of this method crucially depend on the shape of the irradiating continuum, which is likely to be different from the usually assumed power law shape (Zdziarski et al., 2025a), especially in the soft state, when it is apparently formed by Compton upscattering of the disk emission by energetic electrons.
@@
Here, we measure the spin using the continuum method (McClintock et al., 2014) coupled with the reflection method. This is applicable in the soft spectral state. The continuum method relies on the dependence of the innermost stable circular orbit (ISCO) radius on spin (Bardeen et al., 1972), utilizing relativistic models of accretion disks to fit observed X-ray spectra. For high-quality X-ray data, the coupled method can also be used to estimate the mass and distance (Parker et al., 2016), which we follow here.
We study a set of fully simultaneous NICER and NuSTAR observations of this source in the very soft state, i.e., ones with a strong disk blackbody and a very weak high-energy tail. We use two models of relativistic disks, kerrbb (Li et al., 2005) and slimbh (Straub et al., 2011). The former treats modifications of the local disk spectra with respect to the LTE using a color correction, defined as the ratio of the color temperature to the effective one. The latter uses calculations of the vertical radiative transfer of the disk emission (Davis et al., 2005; Davis & Hubeny, 2006), as well as taking into account the increasing disk scale height when the luminosity becomes comparable to the Eddington limit (e.g., Sądowski et al. 2011).
A puzzling feature of this source is the presence of a strong and broad Fe K emission line feature in the presence of a weak high-energy tail, especially in the initial part of the studied observation (Xu et al., 2020). When the tail is fitted by a power law plus reflection, the latter is way too weak to account for that feature. The initial solution proposed by Xu et al. (2020) was that the broad line is intrinsically emitted by the disk, being formed due to the transfer of the blackbody emission from deep in the disk to its surface. However, this disagrees with the standard disk spectra taking into account vertical radiative transfer (Davis et al. 2005; Davis & Hubeny 2006; S. Davis, private communication), which do not show strong Fe K features. Then, Rout et al. (2023) proposed that the Fe K feature is due to the reflection of returning blackbody radiation emitted by the disk (Schnittman & Krolik, 2009). However, Rout et al. (2023) modeled the returning radiation as a single blackbody with a free normalization (García et al., 2022), which can lead to spurious results. Furthermore, they included both the single blackbody and the disk blackbody in their fits, which is clearly incorrect. In the present work, we attempt to resolve this puzzle.
2 The data
The MAXI (Matsuoka et al., 2009) and Swift/BAT (Barthelmy et al., 2005) light curves of the outburst of MAXI J1631–479 are shown in Figure 1 of Xu et al. (2020). We study here the NuSTAR observation denoted as OBS1 in that paper (ObsID 90501301001), which was performed on MJD 58500, after the peak of the MAXI light curve and before the peak of the BAT light curve. Its spectra consist of a strong disk blackbody component and a weak high-energy tail. We define here Part I of the OBS1 as in Figure 1 and Table 1 of Xu et al. (2020). However, we define here Part II excluding a transitional period. In addition, we consider simultaneous observations during the same epochs by NICER, presented in Rout et al. (2021). The log of the observations studied here is given in Table 1, and the light curves of the detectors and the NuSTAR hardness ratio are shown in Figure 1.
The NuSTAR data are used in the 3–79 keV range. They were extracted using the option for bright sources222statusexpr=
"(STATUS==b0000xxx00xxxx000)&&(SHIELD==0)"
. Despite the NuSTAR and NICER being nearly simultaneous, we find the average spectrum from NICER shows some disagreement with the NuSTAR average spectrum in the overlapping energy range (see Section 3). In addition, the NICER spectrum shows a strong positive feature between 1 and 2 keV, which appears to be due to a calibration issue (see, e.g., Hall et al. 2025). Thus, we use the NICER spectrum in the energy range of 2–10 keV. The spectral data have been optimally binned (Kaastra & Bleeker, 2016) with an additional requirement of at least 25 counts per bin.
Table 1: The log of the observations.
@@ 2 The data
3 Spectral fits
3.1 The method
@@
We follow the self-consistent soft-state spectral modeling developed by Zdziarski et al. (2025a). The main new feature of this method is the use of convolution for both calculating the spectrum from Comptonization of the disk emission and calculating the reflection spectrum from the former spectrum irradiating the disk. This avoids the assumption of the incident spectra being power laws with either exponential or thermal-Compton cutoffs, inherent when relxill or relxillCp models (García & Kallman, 2010; García et al., 2018b; Dauser et al., 2016) are used.
@@
As found for Cyg X-1 (Zdziarski et al., 2024b) and GX 339–4 (Zdziarski et al., 2025a), high-energy tails in the soft state require the presence of nonthermal tails beyond the Maxwellian electron distribution of the Comptonizing plasma. The tail used in comppsc is a power law in the electron momentum, and it is given by $(\gamma\beta)^{-p}$ (where $\beta$ is the dimensionless electron velocity) from $\gamma_{\rm min}$, at which the normalizations of the Maxwellian and of the power law are equal, up to $\gamma_{\rm max}$. In our fits, we obtain relatively large values of $p$, for which the fits are independent of $\gamma_{\rm max}$, which we thus set equal to 100.
We use two models of relativistic disks. One is kerrbb (Li et al., 2005), which models geometrically thin disks using the formalism of Novikov & Thorne (1973) and assumes a local spectrum to be a blackbody with a color correction, $f_{\rm col}$, which we assume to equal 1.7 (Davis & El-Abd, 2019). The other is slimbh (described in Straub et al. 2011), based on the calculations of Sądowski (2009, 2011) and Sądowski et al. (2011). It takes into account the finite disk thickness, important for $L\gtrsim 0.1L_{\rm E}$ (Abramowicz et al., 1988). It also treats the local spectra using the radiative-transfer calculations of Davis et al. (2005) and Davis & Hubeny (2006) (when using the option $f_{\rm col}=-1$).
@@
where disk = kerrbb or slimbh. The first term in the curly brackets gives the disk emission alone, and the second term gives the relativistically broadened reflection. Both are subject to Comptonization, given by the comppsc term outside the brackets. The spin parameters in disk and relconv are tied. Then, plabs multiplies the model spectra by $KE^{-\Delta\Gamma}$, where $K=1$ and $\Delta\Gamma=0$ for the NuSTAR A unit, while they are free for the other detectors. The interstellar absorption is modeled by tbfeo (Wilms et al., 2000), where we assume the abundances of Anders & Grevesse (1989), except that those of O and Fe are free (with the relative abundances, $Z$, constrained to 0.5–2.0). We define our photon energy grid by the command energies 0.01 1000 1000 log (necessary for convolution models). We note that increasing the sampling of the energy grid twice (energies 0.01 1000 2000 log) did not influence the fit or its parameters. We use 90% confidence uncertainties ($\Delta\chi^{2}\approx 2.71$; Lampton et al. 1976).
3.2 Individual fits
The spectra of Part I and Part II were initially fitted separately. We first used kerrbb. We obtain good fits, with $\chi_{\nu}^{2}$ values of 531/422 and 638/506, respectively. However, the fitted spins, masses, and distances strongly disagree. In particular, $a_{*}=-0.70_{-0.10}^{+0.32}$, for Part I, while it is positive for Part II, $a_{*}=0.67_{-0.53}^{+0.15}$.
The agreement of the spin parameters is substantially better while using slimbh and the obtained values are positive; $a_{*}=0.85_{-0.07}^{+0.13}$ and $0.81_{-0.10}^{+0.07}$ for Part I and II, respectively. Also, the fits are better, $\chi_{\nu}^{2}=517/422$ and 635/506, respectively. Still, the fitted spins, masses, inclinations, and distances are different. We thus opt for joint fitting as our final approach (Section 3.3).
@@ 3.2 Individual fits
Before that, we show the presence of Fe K residuals and reflection using the individual data. We fit the NuSTAR data alone with a model without reflection, i.e., with Equation 1 without the second term in the curly brackets. We obtain bad fits with $\chi^{2}_{\nu}=1078/297$ and 1012/390, respectively, with strong residuals showing the presence of strong reflection. The data-to-model ratios are shown in Figure 2. We see distinct Fe K features around 7 keV, edges at energies above it, and reflection humps at energies $\gtrsim$10 keV, confirming the result of Xu et al. (2020). However, we see a clear discrepancy between the line profiles from FPM A and B for both data sets, with the latter showing dips around 7 keV. The apparent dip in the NuSTAR FPM B profile was attributed to the presence of an absorption line by Rout et al. (2023). However, this issue is discussed in detail in Draghis et al. (2024), who explain it as due to the proximity of the Sun causing source tracking to be unstable. They attempted to improve the data extraction using a custom procedure. However, the marked difference between the FPM A and B profiles around 7 keV remained, as shown in their Figure 19. Here, we use both detectors, but test for the effect of the removal of the FMP B in Section 3.3 below.
@@
Figure 2: The data to model ratios for the model fitted with slimbh to the NuSTAR data alone and without the reflection component for Part I (left) and Part II (right). We see patterns characteristic of Compton reflection.
3.3 Joint fitting
@@
Figure 3: The NICER (green) and NuSTAR FPM A and B (black and red, respectively) unfolded spectra (top panels) and data-to-model ratios (bottom panels) for the joint fit using slimbh. The Part I and II results are shown in the left and right panels, respectively. The total model spectra and the unabsorbed ones are shown by the solid black and blue curves, respectively. The unscattered parts of the disk emission, scattering alone, and reflection are shown by the magenta, green, and red curves, respectively. Their sums equal the total unabsorbed spectra (in blue).
@@
Disk $L_{\rm disk}/L_{\rm E}$ $0.27^{+0.02}_{-0.02}$ $0.26^{+0.02}_{-0.06}$
Compton $kT_{\rm e}$ [keV] $18.4^{+9}_{-4.5}$ $8.1^{+1.4}_{-1.1}$
$y$ $028^{+0.016}_{-0.008} 0.0$$046^{+0.004}_{-0.007}$
$p$ $3.8_{-0.2}^{+0.2}$ $4.4_{-0.1}^{+0.2}$
$\gamma_{\rm min}$ $29^{+0.07}_{-0.03} 1.1$$12^{+0.02}_{-0.02}$
$f_{\rm cov}$ $0.36^{+0.09}_{-0.13}$ $0.57^{+0.10}_{-0.10}$
@@
$\chi_{\nu}^{2}$ 1169/938
@@
Here, we fit the two data sets jointly. We assume common values of the ISM absorption parameters, mass, distance, inclination, spin, and the Fe abundance of the reflecting medium. We first use kerrbb. We obtain $\chi_{\nu}^{2}=1192/938$ giving a negative spin of $a_{*}=-0.61^{+0.23}_{-0.26}$. However, retrograde accretion is unlikely for an LMXB, see also the discussion of past claims of negative spins in LMXBs in Zdziarski et al. (2025a).
@@
Given that, we decide on slimbh to be our final model, which also yields a better fit with $\chi_{\nu}^{2}=1169/938$, giving a relatively high and positive spin of $a_{*}=0.84\pm 0.03$; see Table 2 for all of the parameters. The corona covers roughly half of the disk, with $f_{\rm cov}\sim 0.5$. The obtained mass is within the typical range for BHs in LMXBs (Özel et al., 2010), and the inclination agrees with that of Xu et al. (2020). We note that slimbh does not allow for $a_{*}<0$. However, the $\chi^{2}$ dependence on the decreasing $a_{*}$ is approximately monotonically increasing, with $\Delta\chi^{2}=+34$ at $a_{*}=0$. Thus, the presence of a local minimum at $a_{*}<0$ appears unlikely. The spectra and residuals are shown in Figure 3. An important feature of these spectra is that the scattered disk emission, which is both emitted outside and impinges on the disk, has a shape very different from a power law. We see no significant residuals for the NuSTAR data (except for the highest energies), but we see that the NICER spectrum shows wavy residuals, indicative of instrumental issues. The apparent line above 2 keV is due to the gold edge of the NICER detector. Since its contribution to $\chi^{2}$ is small, we opted not to model it.
Then, we study the effect of the removal of the FMP B data (which can be affected by the proximity of the Sun, see Section 2). We found its effect on the fitting parameters to be minor, with most of the best-fit parameters having almost the same values, while the uncertainty ranges increase. The spin parameter becomes $a_{*}=0.83^{+0.05}_{-0.08}$. The largest difference is in the value of the reflection strength for Part I, which is ${\cal R}=1.8{+1.2}_{-0.7}$, extending to lower values than the full data, ${\cal R}=2.4^{+0.6}_{-0.4}$.
@@
Finally, we consider a model with a warm and optically thick corona on top of the accretion disk, considered before for Cyg X-1, LMC X-1, M33 X-7, and GX 339–4 (Belczyński et al., 2024; Zdziarski et al., 2024a, b, 2025a). Such coronae were modeled there by optically thick thermal Comptonization using the code thcomp (Zdziarski et al., 2020). The differences of this model with respect to that of Equation 1 is the addition of the thcomp factor in front of the disk factor and that a free color correction in slimbh is used instead of the model including radiative transfer (see a discussion in Zdziarski et al. 2025a). However, we have not been able to find a model with a warm corona that would provide a better fit than our main model. The best one had $\chi_{\nu}^{2}=1180/933$. Thus, we conclude that such a component may not be present in MAXI J1631–479.
4 Discussion

### v2

@@ Abstract
We study the transient black hole binary MAXI J1631–479 observed simultaneously by NICER and NuSTAR in its soft spectral state. Its puzzling feature is the presence of a strong and broad Fe K line, while the continuum includes a prominent disk blackbody and a very weak power-law tail. The irradiation of the disk by a power-law spectrum fitting the tail is far too weak to explain the strong line. Previous proposals included the idea that the Fe K emission is intrinsic to the disk. Here, we propose that the strong line can be explained by the irradiation of the disk by photons from Comptonization of the disk blackbody by coronal electrons. One crucial effect is that the shape of the irradiating spectrum at $\lesssim$10 keV reflects that of the disk blackbody; it is strongly curved and has a higher flux than what would be produced by a fit with a power-law irradiation. The other effect is a relativistic enhancement of the backscattered coronal flux incident on the disk. Both effects together can account for the line, although the latter is modeled only quantitatively. While this result is independent of the physical model used for disk emission, the fitted spin depends heavily on that model. When employing a Kerr disk model for a thin disk with color correction, the fitted spin appears retrograde, rare for a Roche-lobe overflow binary. A model that accounts for both the finite thickness of the disk and radiative transfer yields a spin of $a_{*}\approx 0.8$–0.9.
@@ I Introduction
Its BH mass, $M$, and the distance, $D$, remain uncertain. Rout et al. (2023) claimed a very high BH mass, $26{M}_{\sun}$. This was implied by the disk blackbody model seen at high inclination, $i\approx 71\pm 1\arcdeg$, found by them. On the other hand, low inclination values were found in the spectral studies by Xu et al. (2020), $i=29\pm 3\arcdeg$, and by Draghis et al. (2024), $i=22^{+10}_{-12}\arcdeg$. Its dimensionless spin parameter was estimated to be very high, $a_{*}>0.94$ (Xu et al., 2020), $a_{*}>0.996$ (Rout et al., 2023), and $a_{*}=0.95^{+0.04}_{-0.08}$ (Draghis et al., 2024). Those studies used the X-ray reflection method (Bambi et al., 2021) in the soft spectral states of MAXI J1631–479. However, the results of this method crucially depend on the shape of the irradiating continuum, which is likely to be different from the usually assumed power law shape (Zdziarski et al., 2025), especially in the soft state, when it is apparently formed by Compton upscattering of the disk emission by energetic electrons.
@@
Here, we measure the spin using the continuum method (McClintock et al., 2014) jointly with the reflection method, which is applicable in the soft spectral state (Parker et al., 2016). The continuum method relies on the dependence of the innermost stable circular orbit (ISCO) radius on spin (Bardeen et al., 1972) and uses relativistic accretion disk models to fit observed X-ray spectra. We study a set of simultaneous NICER and NuSTAR observations of this source in the very soft state, i.e., ones with a strong disk blackbody and a very weak high-energy tail. We use two models of relativistic disks, kerrbb (Li et al., 2005) and slimbh (Straub et al., 2011).
Table 1: The log of the observations
@@
A puzzling feature of this source is the presence of a strong and broad Fe K emission line feature, with EW $\approx 180$–210 eV (Xu et al., 2020), in the presence of a weak high-energy tail. When the tail is fitted by a power law plus reflection, the latter is way too weak to account for that feature, especially in the initial part of the studied observation (Xu et al., 2020). The initial solution proposed by Xu et al. (2020) was that the broad line is intrinsically emitted by the disk, being formed due to the transfer of the blackbody emission from deep in the disk to its surface. However, this contradicts standard disk spectra that account for vertical radiative transfer (Davis et al. 2005; Davis & Hubeny 2006; S. Davis, private communication), which do not exhibit strong Fe K features. On the other hand, there is an ongoing debate on the applicability of the standard disk model to the soft state of BH XRBs (which we discuss in Section IV), with the modified models still lacking detailed spectral predictions. Then, the Fe K line might still be intrinsic to the accretion disk in this source, as proposed by Xu et al. (2020).
Later, Rout et al. (2023) proposed that the Fe K feature is due to the reflection of blackbody radiation emitted by the disk and returning to it in strong gravity (Cunningham, 1975; Schnittman & Krolik, 2009). However, the model of Rout et al. (2023) included the disk blackbody, the irradiating blackbody (dominant), and its reflection in their fits. However, the irradiating blackbody should not be included at all as a separate component since it is already accounted for by the disk blackbody.
In the present work, we propose a new solution, in which the Fe K feature is still due to the reflection of the coronal radiation due to Compton scattering of the disk blackbody by relativistic electrons. That reflection can be significantly stronger than that from a power-law spectrum fitted to the high-energy tail due to both the strong concave curvature of the actual spectrum from Comptonization, and the relativistic enhancement of the emission toward the disk with respect to that outside (Ghisellini et al., 1991).
II The data and variability
The MAXI (Matsuoka et al., 2009) and Swift/BAT (Barthelmy et al., 2005) light curves of the outburst are shown in Figure 1 of Xu et al. (2020). We study here the NuSTAR observation denoted as OBS1 in that paper (ObsID 90501301001), which was performed on MJD 58500, after the peak of the MAXI light curve and before the peak of the BAT light curve. Its spectra consist of a strong disk blackbody component and a weak high-energy tail. We define here Part I of the OBS1 as in Figure 1 and Table 1 of Xu et al. (2020). However, we define here Part II excluding a short transitional period after Part I (between the vertical dashed curves in Figure 1). In addition, we consider simultaneous NICER observations within the same epochs, as presented in Rout et al. (2021). The log of the observations studied here is given in Table 1, and the light curves of the detectors and the NuSTAR hardness ratio are shown in Figure 1.
The NuSTAR data are used in the 3–79 keV range. We processed them using the standard task nupipeline in the NuSTARDAS package, using the option for bright sources222statusexpr=
"(STATUS==b0000xxx00xxxx000)&&(SHIELD==0)"
. Source and background spectra were extracted using the task nuproducts from circular regions of $100\arcsec$ radius, centered on the source position and from a source-free region, respectively. The corresponding response matrix files and auxiliary response files were generated simultaneously. Despite the NuSTAR and NICER exposures being nearly simultaneous, we find the average spectrum from NICER shows some disagreement with the NuSTAR average spectrum in the overlapping energy range (see Section III). During the reduction pipeline of NICER, the script nicerl3-spect added 1.5% systematic error [^3] https://heasarc.gsfc.nasa.gov/lheasoft/ftools/headas/nicerl3-spect.html to the detector channels in the 0.5–9.0 keV range, and more (up to 2.5%) above 9 keV. In addition, the NICER spectrum shows a strong positive feature between 1 and 2 keV, which appears to be due to a calibration issue (see, e.g., Hall et al. 2025). Thus, we use the NICER spectrum in the 2–10 keV energy range. The spectral data have been optimally binned (Kaastra & Bleeker, 2016), with the additional requirement that each bin contain at least 25 counts.
We have also calculated the fractional variability of the light curves, using both the NICER and NuSTAR data. The results for the frequency range of 0.01–1 Hz are shown in Table 2. We see that the data up to 15 keV show very weak variability. The accretion disk spectrum dominates this range. On the other hand, the high-energy tail exhibits strong variability, particularly in Part I.
III Spectral fits
III.1 The method
@@
We follow the self-consistent soft-state spectral modeling developed by Zdziarski et al. (2025). The main new feature of this method is the use of convolution for both calculating the spectrum of Comptonized disk emission and the reflection spectrum from the former spectrum irradiating the disk. This avoids the assumption that the incident spectra are power laws with either exponential or thermal Compton cutoffs, inherent in either relxill or relxillCp models (García & Kallman, 2010; García et al., 2018b; Dauser et al., 2016).
Table 2: The fractional rms variablity in the 0.01–1 Hz frequency range
Detector Energy range rms Part I rms Part II
NICER 2–10 keV $0.65\pm 0.08\%$ $0.62\pm 0.06\%$
NuSTAR 3-15 keV $0.64\pm 0.05\%$ $0.60\pm 0.06\%$
NuSTAR 15-25 keV $16.4\pm 3.3\%$ $4.4\pm 3.4\%$
@@
High-energy tails in the soft state, when modeled by Comptonization, require the presence of nonthermal tails beyond the Maxwellian electron distribution, in particular in Cyg X-1 (Zdziarski et al., 2024b) and GX 339–4 (Zdziarski et al., 2025). The tail used in comppsc is a power law in the electron momentum, i.e., $(\gamma\beta)^{-p}$ (where $\beta$ is the dimensionless electron velocity) from $\gamma_{\rm min}$, at which the normalizations of the Maxwellian and of the power law are equal, up to $\gamma_{\rm max}$. In our fits, we obtain relatively large values of $p$, for which the fits are independent of $\gamma_{\rm max}$, which we thus set equal to 100.
We use two models of relativistic disks. One is kerrbb (Li et al., 2005), which models geometrically thin disks using the formalism of Novikov & Thorne (1973) and assumes a local spectrum to be a blackbody with a color correction, $f_{\rm col}$, which we assume to equal 1.7 (Davis & El-Abd, 2019). The other is slimbh (described in Straub et al. 2011), based on the calculations of Sądowski (2009, 2011) and Sądowski et al. (2011). It accounts for the finite disk thickness, which is important for $L\gtrsim 0.1L_{\rm E}$ (Abramowicz et al., 1988). Furthermore, it treats the local spectra using the radiative-transfer calculations of Davis et al. (2005) and Davis & Hubeny (2006) when the option $f_{\rm col}=-1$ is set. These spectra depend on the viscosity parameter, which we set as $\alpha=0.1$.
@@
where disk = kerrbb or slimbh. The first term in the curly brackets gives the disk emission alone, and the second term gives the relativistically broadened reflection. Both are subject to Comptonization, given by the comppsc term outside the brackets.
Then, plabs accounts for small differences in the spectral slope between the NuSTAR A and B units and the NICER, which we found clearly present in the data. We note that the standard xspec model plabs is equivalent to the external model CRABCORR, defined by Steiner et al. (2010) for this purpose and widely used. Both multiply the model spectra by $KE^{-\Delta\Gamma}$, where $K=1$ and $\Delta\Gamma=0$ for the NuSTAR A unit, while they are free for NuSTAR B and NICER. We found that requiring $\Delta\Gamma=0$ for NuSTAR B increases the $\chi^{2}$ of the fits by $\sim$120. The interstellar absorption is modeled by tbfeo (Wilms et al., 2000), where we assume the abundances of Wilms et al. (2000). We have found that while the O relative abundance is compatible with unity (which we hereafter set), allowing the Fe abundance to be free reduces $\chi^{2}$ by $\sim$10. Convolution models require the energy grid to be substantially larger (on both sides) than the detector’s grid, and we thus set it using the command energies 0.01 1000 2000 log. We use 90% confidence uncertainties ($\Delta\chi^{2}\approx 2.71$; Lampton et al. 1976).
III.2 Individual fits
@@
We initially fitted the spectra of Parts I and II separately. We first demonstrate the presence of Fe K residuals and reflection using individual data. We fit the NuSTAR data alone without reflection with slimbh, i.e., using Equation 1 without the second term in the curly brackets. We obtain poor fits with $\chi^{2}_{\nu}=1120/298$ and 1087/391, respectively, with strong residuals indicating significant reflection. The data-to-model ratios are shown in Figure 2. We see distinct Fe K features, edges at energies above them, and reflection humps at $E\gtrsim 10$ keV, confirming the result of Xu et al. (2020). However, we observe a clear discrepancy in both Parts I and II of the profiles in the 7.00–7.37 keV range between the A and B units, with the latter showing clear dips. These dips were attributed to the presence of an absorption line by Rout et al. (2023). On the other hand, this issue was discussed in detail in Draghis et al. (2024), who attributed it to the Sun’s proximity, which destabilizes source tracking, and to NUSTAR B being partly in the detector gap. They attempted to improve data extraction by implementing a custom procedure. However, the marked difference between the A and B profiles around 7 keV remained virtually unchanged, as shown in their Figure 19. Therefore, we kept using the standard data extraction, but we excluded the 7.00–7.37 keV from the NuSTAR B. In that case, we find full agreement between the two units (after applying the small-slope correction; see Table 3).
We fit the NICER+NuSTAR data with slimbh, including the reflection. We obtain $\chi_{\nu}^{2}$ values of 493/419 and 602/503, respectively. We find well-constrained and similar values of the spin parameter, $a_{*}=0.88_{-0.03}^{+0.06}$ and $0.83_{-0.06}^{+0.04}$ for Part I and II, respectively. The fitted spins, masses, inclinations, and distances are also relatively similar.
We then use kerrbb with reflection. We obtain $\chi_{\nu}^{2}$ values of 511/419 (significantly higher than that for the model with slimbh) and 604/503 for Parts I and II, respectively. The spin is only weakly constrained, $a_{*}=-0.60_{-0.24}^{+0.33}$ and $0.26_{-0.74}^{+1.02}$ for Parts I and II, respectively.
@@
Figure 2: The data to model ratios for the model fitted with slimbh to the NuSTAR data alone and without the reflection component for Part I (left) and Part II (right). We observe patterns characteristic of Compton reflection, as well as a marked difference in the spectra from the two units around 7 keV.
III.3 Joint fitting
@@
Figure 3: The NICER (green) and NuSTAR A and B (black and red, respectively) unfolded spectra (top panels) and data-to-model ratios (bottom panels) for the joint fit using slimbh. The Part I and II results are shown in the left and right panels, respectively. The total model spectra and the unabsorbed ones are shown by the solid black and blue curves, respectively. The unscattered disk emission, scattering alone, and reflection are shown by the magenta, green, and red curves, respectively. Their sums equal the total unabsorbed spectra (in blue).
@@
Disk $L_{\rm disk}/L_{\rm E}$ $0.26^{+0.01}_{-0.01}$ $0.25^{+0.03}_{-0.03}$
Compton $kT_{\rm e}$ [keV] $18^{+4}_{-4}$ $8.2^{+1.9}_{-1.0}$
$\tau$ $0.25^{+0.30}_{-0.10}$ $0.68^{+0.29}_{-0.18}$
$p$ $3.7_{-0.1}^{+0.2}$ $4.4_{-0.1}^{+0.2}$
$\gamma_{\rm min}$ $29^{+0.08}_{-0.02} 1.1$$12^{+0.02}_{-0.03}$
$f_{\rm cov}$ $0.28^{+0.23}_{-0.10}$ $0.61^{+0.12}_{-0.12}$
@@
Cross-calibration $\Delta\Gamma_{\rm NICER}$ $0.11_{-0.01}^{+0.01}$
$K_{\rm NICER}$ $0.84_{-0.01}^{+0.01}$ $0.85_{-0.01}^{+0.01}$
$\Delta\Gamma_{\rm NuSTAR\,B}$ $0.017_{-0.002}^{+0.003}$
$K_{\rm NuSTAR\,B}$ $2_{-0.01}^{+0.01} 1.010$$1_{-0.01}^{+0.01}$
$\chi_{\nu}^{2}$ 1108/931
@@
Here, we jointly fit the two data sets. We assume common values of the interstellar absorption parameters, mass, distance, inclination, spin, and the Fe abundance of the reflecting medium. We first use kerrbb. We obtain $\chi_{\nu}^{2}=1129/931$ giving a negative spin of $a_{*}=-0.40^{+0.15}_{-0.23}$. We have searched the parameter space for a solution with $a_{*}>0$, but none has been found.
Then, we use slimbh, which yields a better fit with $\chi_{\nu}^{2}=1108/930$. It gives a well-constrained, positive, and relatively high spin of $a_{*}=0.86^{+0.03}_{-0.03}$. It is also a more physically motivated model, incorporating the dependence on the Eddington ratio through both the disk scale height and the radiative transfer effects (Straub et al., 2011). We then selected slimbh as our final model. See Table 3 for all the parameters. The spectra and residuals from the joint fit are shown separately for Parts I and II in Figure 3. We have also tested the effect of not including the correction for the spectral slope between the NuSTAR A and B units. When setting $\Delta\Gamma_{\rm NuSTAR\,B}=0$, we obtain a significantly worse fit, $\Delta\chi^{2}=+119$, but the spin remains the same with a slightly lower uncertainty, $a_{*}=0.86^{+0.03}_{-0.02}$.
@@
The corona covers roughly half of the disk. The obtained mass is within the range found for BHs in LMXBs (Özel et al., 2010). We note that slimbh does not allow for $a_{*}<0$. However, the $\chi^{2}$ dependence on the decreasing $a_{*}$ is approximately monotonically increasing, and $\Delta\chi^{2}=+45$ at $a_{*}=0$. Thus, the presence of a local minimum at $a_{*}<0$ appears unlikely. An essential feature of the spectra shown in Figure 3 is that the scattered disk emission (green), which is both emitted outside and impinges on the disk, has a shape very different from a power law. We observe no significant residuals in the NuSTAR data, but the NICER spectrum shows wavy residuals, indicative of instrumental issues. The apparent line above 2 keV is due to the gold edge of the NICER detector. Since its contribution to $\chi^{2}$ is small, we opted not to model it.
@@
Finally, we consider a model with a warm, optically thick corona on top of the accretion disk, as previously applied to Cyg X-1, LMC X-1, M33 X-7, and GX 339–4 (Belczyński et al., 2024; Zdziarski et al., 2024a, b, 2025). In those cases, the presence of such coronae allowed spin values close to null, and significantly improved the fit for Cyg X-1. Such coronae were modeled there by optically thick thermal Comptonization using the code thcomp (Zdziarski et al., 2020). The differences of this model with respect to that of Equation 1 is the addition of the thcomp factor in front of the disk factor and that a free color correction in slimbh is used instead of the model including radiative transfer (see a discussion in Zdziarski et al. 2025). However, we have not identified a model with a warm corona that yields a better fit than our main model. The best fit has $a_{*}=0.00^{+0.08}$ and $\chi_{\nu}^{2}=1120/926$. Thus, we conclude that while such a component may still be present in MAXI J1631–479, there is no statistical evidence for it.
IV Discussion